# Project Riley: Multimodal Multi-Agent LLM Collaboration with Emotional Reasoning and Voting

### v1

@@ Abstract
Preprint Notice: This manuscript is a preprint and is currently under review for potential publication in the journal Information Fusion.
@@ 1 Introduction
The ‘Inside Out’ saga [1, 2], produced by Pixar Animation Studios, provides a symbolic narrative depiction of the perceptions and emotional interactions inherent in human experience.
Throughout the development of the saga’s films, the creators aimed not only to produce artistic entertainment but also to translate fundamental psychological concepts into a language accessible to the general public. Consequently, the films benefited from consultation with experts in psychology and neuroscience, including Dacher Keltner, Paul Ekman, and Lisa Damour [3, 4].
The saga narrates the life of a girl named Riley, with the use of autonomous characters representing distinct emotions serving to personify her emotional experiences. The first film (2015) [1], is set during Riley’s childhood and focuses on five primary emotions: Joy, Sadness, Fear, Anger, and Disgust. The sequel (2024) [2] portrays Riley’s adolescence and the complexities typical of this developmental stage, introducing additional emotions: Shame; Envy; Boredom; and Anxiety. These emotions interact dynamically, guiding Riley’s behavioural responses to life changes and challenges.
@@
Therefore, this paper presents the architecture of the generative AI conversational solution, Project Riley. This system is characterised by its emotional awareness, multi model functionality, and the implementation of sophisticated self-refining and voting logic. Project Riley integrates multimodal information—text and images—through the coordinated use of multiple Large Language Models (LLM), each representing a distinct emotion. It processes user input using both standard and multimodal LLM for textual and visual analysis, respectively, and orchestrate multi-round emotional dialogues. The final output results from a structured voting and synthesis mechanism that uses an advanced reasoning model to combine emotionally differentiated responses into a coherent answer. This architecture exemplifies the potential of GenAI for symbolic and neural fusion, by coupling the generative capabilities of neural models with symbolic reasoning structures—such as agent-specific roles, conversational history tracking, voting with justification, and formally segmented outputs—thereby enabling interpretation and decision-making across heterogeneous data sources.
@@
To the best of our knowledge, this is the first framework that leverages generative AI to orchestrate structured affective reasoning through independent emotional agents, introducing a novel paradigm for information fusion via generative emotional modelling. Furthermore, no comparable solution has been identified in the literature that adapts such an architecture to emergency response contexts as implemented in Armando, where emotional regulation, informational accuracy, and user trust are prioritised under high-stakes conditions.
@@ 2.1 Emotion Recognition and Emotional Intelligence
Recent literature on emotionally intelligent systems primarily addresses the integration of emotional awareness into AI models, particularly LLM. Ratican et al. [5] introduced the 6DE conceptual model, featuring six emotional dimensions—arousal, valence, dominance, agency, fidelity, and novelty—to provide a comprehensive framework for analysing human emotions in AI systems. This approach aims to enhance empathy and contextual relevance, specifically targeting applications in education, mental health, and assisted care. Their suggested integration methods include emotion-guided prompting, annotated datasets, user feedback, emotional database integration, and emotional quality refinement through post-processing.
@@
Kang et al. [6] similarly focused on emotional integration, albeit within a social robotic context, employing the Pleasure-Arousal-Dominance (PAD) emotional model. They developed the SoR-ReAct agent for the Nadine robot platform, which combines GPT-4 with personalised episodic memory, enabling emotional simulation and personalised interactions based on historical data. Unlike Ratican et al., Kang et al. directly incorporated episodic memory retrieval facilitated by OpenAI’s embeddings, enhancing the robot’s interactive empathy capabilities through multimodal perception.
@@
In contrast, Liu et al. [7] addressed emotional intelligence specifically within negotiation scenarios, developing EQ-Negotiator, which integrates emotion recognition (RoBERTa, GPT-2, DeBERTa) with strategic game-theory-based reasoning. Their work distinctly emphasised balancing emotional empathy and assertiveness, using Hidden Markov Models for dynamic emotional adaptations. Notably, their approach outperformed GPT-4 in terms of emotional realism and negotiation efficacy.
@@
Rasool et al. [8] and Liu et al. [9] extended emotional intelligence into psychotherapy and multimodal dialogue, respectively. Rasool et al. employed emotional lexicons (NRC Emotion Lexicon, SentiWordNet) combined with hierarchical segmentation and attention mechanisms to enhance empathetic responses, despite noting a trade-off between empathy and coherence. Liu et al., conversely, integrated emotional and visual modalities using Emotional Retrieval Module(ERM), Response Emotion Prediction (REP), and Emotion-Enhanced Response Generation (EERG) models. Their multimodal ELMD system demonstrated superior emotional and contextual responsiveness compared to traditional dialogue systems.
@@
Chen et al. [10], Motwani et al. [11], and Yang et al. [12] further highlighted emotional considerations, albeit within specialised frameworks. Chen et al.’s RECONCILE framework collaboratively utilised multiple LLM agents to generate emotionally and contextually relevant responses, leveraging weighted voting based on reasoning justification. Motwani et al. proposed a sequential training method employing specialised agents in reasoning and verification tasks, indirectly enhancing emotional intelligence through robust error correction. Yang et al.’s SupportlyChat directly focused on therapeutic contexts using Cognitive Behavioural Therapy (CBT), combining RoBERTa and sentiment analysis to enhance emotional awareness and professional interaction in mental health support scenarios.
@@
Finally, Brun et al. [13] evaluated emotional sensitivity in technical support chatbots, specifically employing VADER sentiment analysis within ChatGPT-3.5. Their findings indicated emotional sensitivity significantly improved user-perceived competence, empathy, trust, and continued use intention, although without measurably affecting user emotional states.
@@ 2.2 Multi-Agent Debate and Collaborative Reasoning
Research exploring multi-agent systems frequently emphasises collaborative reasoning and decision-making through debate mechanisms. Zhao et al. [14] tested electoral methods from social choice theory (e.g., Borda, IRV, Ranked Pairs) in collective decision-making scenarios involving multiple LLM agents, highlighting notable accuracy improvements in benchmarks like MMLU and ARC. They demonstrated that even minimal agent ensembles significantly enhance performance and recommended task-specific tailoring of decision methods.
Yang et al. [15] combined adversarial debate, weighted voting, and internal self-correction within a multi-agent framework to mitigate hallucinations and enhance reasoning efficiency. Their methodology involved iterative debate, error documentation, and dynamic weighting, outperforming baseline methods in accuracy and response speed on tasks like GSM8K and MMLU.
Similarly, Xu et al. [16] proposed a peer-review inspired collaborative method. Agents independently generated solutions, reviewed peers’ contributions, and refined responses accordingly. This explicit feedback and diverse model integration consistently improved performance across multiple datasets (GSM8K, StrategyQA, ARC-c).
@@
In contrast to existing approaches, the architecture proposed in this work distinguishes itself by employing a structured multi-agent framework in which each agent represents a distinct emotion and participates in iterative emotional reasoning. While prior studies tend to integrate emotion either through static embeddings, sentiment lexicons, or personalised memory modules [5, 6, 8], this architecture fosters dynamic emotional dialogue and collective deliberation. Additionally, the integration of symbolic reasoning mechanisms—such as voting and justification—enhances explainability, a feature largely absent in previous neural architectures [10, 11]. The system is also extensible to different emotional models and has been adapted for high-stakes domains, such as emergency response, through the Armando variant. To the best of our knowledge, no previous work combines emotional simulation, multimodal reasoning, and factual grounding within a unified architecture.
@@ 3 Proposed Architecture
This paper presents an innovative conversational AI system architecture that processes user queries through distinct emotional lenses, specifically a set of five basic emotions: Anger, Joy, Sadness, Fear, and Disgust. The system follows the emotional framework portrayed in the first Inside Out film, which was inspired by Paul Ekman’s theory of universal emotions [17]. Although Ekman initially identified six basic emotions—Happiness, Anger, Sadness, Disgust, Surprise, and Fear—and later expanded the list to include emotions such as Contempt and Enjoyment [18], the film deliberately reduced the set to five for narrative clarity [19, 20]. In this work, the same five emotions were retained to preserve the analogy with the film, thereby fostering greater familiarity and intuitive use—particularly among users who have seen the film. However, it is important to emphasise that the Project Riley architecture is not limited to this configuration: it is capable of simulating a broader range of emotions, depending on the requirements of the application and the computational resources available.
The system can be easily adapted to reflect alternative emotional models, allowing developers to incorporate different sets of emotions based on specific application needs or cultural relevance. Furthermore, the proposed architecture can be aligned with other psychological theories, such as Plutchik’s Wheel of Emotions [21], Carroll Izard’s Differential Emotions Theory [22], Jaak Panksepp’s Affective Neuroscience framework [23], or dimensional approaches like Russell’s Circumplex Model [24] and the Pleasure-Arousal-Dominance (PAD) model [25]. Such adaptations would require architectural modifications and prompt reengineering to ensure consistency with the theoretical assumptions of the chosen model.
@@ 4 Project Riley chatbot prototype
In accordance with the proposed architecture, a general-purpose chatbot was implemented based entirely on the Project Riley architecture. To enable local deployment of LLM, the Ollama framework was utilised [26]. The hardware specifications of the machine used for this implementation are detailed in Table 1.
@@
Given the hardware constraints, particularly regarding Graphics Processing Unit (GPU) Video Random Access Memory (VRAM) capacity, considerable attention was directed towards selecting models that offer a balanced trade-off between performance and model size. Furthermore, when computationally feasible, we prioritised the use of abliterated models. Abliteration is a technique that alters the weights of a model to bypass default safety alignment mechanisms typically introduced during fine-tuning [27]. These mechanisms, while intended to ensure safe output, often over-filter emotionally charged or controversial prompts, leading to overly polite, neutral, or evasive responses.
@@
Consequently, the chosen text-based LLM model was huihui_ai/llama3.2-abliterate:3b [28], which has 3.61 billion parameters and a model size of 2.2 GB. For image description tasks, the selected LLM was gemma3:4b [29], a model capable of processing both text and images, comprising 4.3 billion parameters and a model size of 3.3 GB. For advanced reasoning tasks, the huihui_ai/deepseek-r1-abliterated:8b [30] model was utilised, containing 8.03 billion parameters and a model size of 4.9 GB.
@@ 4.1 Prototype Evaluation
Analysis of responses to the question Which emotion do you predominantly identify in the final response? (Table 3) reveals significant variability in emotional perception. In the Anxiety in Academic or Professional Contexts theme, Joy was most frequently identified (7 mentions), possibly reflecting a reassuring approach adopted by the system. Conversely, Sadness dominated responses related to Breakups or Friendship Loss (5 mentions), indicating the appropriateness of a more reflective emotional tone in relational loss scenarios. Difficult Personal Decisions equally elicited Fear and Joy (5 mentions each), suggesting responses were balanced between cautionary and optimistic perspectives. Family or Intergenerational Conflicts predominantly evoked Joy (6 mentions), again highlighting positivity in familial contexts. Lastly, Job Loss or Unemployment mainly elicited Fear (5 mentions), aligning well with the typically uncertain and stressful nature of employment-related concerns.
@@ 5 Armando: A emergency response chatbot using Project Riley
This approach seeks to mitigate issues such as those observed on 28 April 2025, during a widespread power outage across the Iberian Peninsula, when some individuals turned to ChatGPT for information. As the chatbot was not updated with official communications from public safety authorities, its use led to the dissemination of unreliable and inaccurate information, thereby increasing public confusion and anxiety [31].
To ensure that Armando’s responses accurately reflect official guidelines and authoritative information, a Retrieval-Augmented Generation (RAG) approach has been implemented. RAG enhances chatbot interactions by integrating external knowledge-context injection-during chatbot response generation. Specifically, a diverse set of authoritative sources—comprising official documents and information extracted from verified webpages—is indexed through semantic embeddings. These embeddings, generated using the mxbai-embed-large model in this case, provide a structured representation of the underlying content in a high-dimensional vector space. When a user submits a query, its semantic embedding is computed and compared against the indexed document embeddings. The content with the highest semantic similarity is then retrieved and used to ground the system’s response in accurate and contextually relevant information. When a user interacts with Armando, the chatbot retrieves relevant information associated with embeddings that closely match the query context, enriching its generated responses with verified external knowledge.
@@
Given the specific requirements of this emergency response application, namely, reduced response times and lower computational overhead—targeted optimisations were implemented. The Final Synthesis and Voting and Analysis processes were simplified to enhance execution speed and reduce GPU VRAM usage. To this end, prompt modifications were introduced to emulate a reasoning logic using a conventional text-based LLM. Furthermore, as the THOUGHTS section was deemed non-essential in an emergency context and its exclusion contributed to faster response generation, it was omitted from the chatbot’s Final Synthesis.
@@
As illustrated in Figure 5, RAG enrichment is applied immediately before the final synthesis stage—contrary to what might appear more intuitive, which would be to introduce it prior to the multi-round processing. Such an approach would provide each emotional agent with validated information from the outset, allowing the entire discussion to be grounded in authoritative content. However, initial development and empirical testing revealed that this strategy does not reliably ensure the factual accuracy of the final output, as critical information may become diluted or distorted during the emotional dialogue. Therefore, RAG enrichment was instead positioned just before the synthesis stage, ensuring that the final response maintains emotional balance and contextual sensitivity, while also conveying accurate and up-to-date information to the user.

### v2

@@ Abstract
keywords:
LLM, emotionally-aware self-refining, conversational generative AI, RAG, emergency response
@@ 1 Introduction
The ‘Inside Out’ saga Docter2015 ; Mann2024 , produced by Pixar Animation Studios, provides a symbolic narrative depiction of the perceptions and emotional interactions inherent in human experience.
Throughout the development of the saga’s films, the creators aimed not only to produce artistic entertainment but also to translate fundamental psychological concepts into a language accessible to the general public. Consequently, the films benefited from consultation with experts in psychology and neuroscience, including Dacher Keltner, Paul Ekman, and Lisa Damour Brice2015 ; Barber2024 .
The saga narrates the life of a girl named Riley, with the use of autonomous characters representing distinct emotions serving to personify her emotional experiences. The first film (2015) Docter2015 , is set during Riley’s childhood and focuses on five primary emotions: Joy, Sadness, Fear, Anger, and Disgust. The sequel (2024) Mann2024 portrays Riley’s adolescence and the complexities typical of this developmental stage, introducing additional emotions: Shame; Envy; Boredom; and Anxiety. These emotions interact dynamically, guiding Riley’s behavioural responses to life changes and challenges.
@@
Therefore, this paper presents the architecture of the generative AI conversational solution, Project Riley. This system is characterised by its emotional awareness, multi model functionality, and the implementation of sophisticated self-refining and voting logic. Project Riley integrates multimodal information, text and images, through the coordinated use of multiple Large Language Models (LLM), each representing a distinct emotion. It processes user input using both standard and multimodal LLM for textual and visual analysis, respectively, and orchestrate multi-round emotional dialogues. The final output results from a structured voting and synthesis mechanism that uses an advanced reasoning model to combine emotionally differentiated responses into a coherent answer. This architecture demonstrates the potential of GenAI to combine the generative capabilities of neural models with symbolic reasoning structures, such as agent-specific roles, conversational history tracking, voting with justification, and formally segmented outputs, thereby enabling interpretation and decision-making across heterogeneous data sources.
@@
To the best of our knowledge, this is the first framework that leverages generative AI to orchestrate structured affective reasoning through independent emotional agents, introducing a novel paradigm for integrating generative emotional modelling into conversational systems. Furthermore, no comparable solution has been identified in the literature that adapts such an architecture to emergency response contexts as implemented in Armando, where emotional regulation, informational accuracy, and user trust are prioritised under high-stakes conditions.
@@ 2.1 Emotion Recognition and Emotional Intelligence
Recent literature on emotionally intelligent systems primarily addresses the integration of emotional awareness into AI models, particularly LLM. Ratican et al. Ratican2023 introduced the 6DE conceptual model, featuring six emotional dimensions — arousal, valence, dominance, agency, fidelity, and novelty — to provide a comprehensive framework for analysing human emotions in AI systems. This approach aims to enhance empathy and contextual relevance, specifically targeting applications in education, mental health, and assisted care. Their suggested integration methods include emotion-guided prompting, annotated datasets, user feedback, emotional database integration, and emotional quality refinement through post-processing.
@@
Kang et al. Kang2024 similarly focused on emotional integration, albeit within a social robotic context, employing the Pleasure-Arousal-Dominance (PAD) emotional model. They developed the SoR-ReAct agent for the Nadine robot platform, which combines GPT-4 with personalised episodic memory, enabling emotional simulation and personalised interactions based on historical data. Unlike Ratican et al., Kang et al. directly incorporated episodic memory retrieval facilitated by OpenAI’s embeddings, enhancing the robot’s interactive empathy capabilities through multimodal perception.
@@
In contrast, Liu et al. Liu2025 addressed emotional intelligence specifically within negotiation scenarios, developing EQ-Negotiator, which integrates emotion recognition (RoBERTa, GPT-2, DeBERTa) with strategic game-theory-based reasoning. Their work distinctly emphasised balancing emotional empathy and assertiveness, using Hidden Markov Models for dynamic emotional adaptations. Notably, their approach outperformed GPT-4 in terms of emotional realism and negotiation efficacy.
@@
Rasool et al. Rasool2024 and Liu et al. Liu2024 extended emotional intelligence into psychotherapy and multimodal dialogue, respectively. Rasool et al. employed emotional lexicons (NRC Emotion Lexicon, SentiWordNet) combined with hierarchical segmentation and attention mechanisms to enhance empathetic responses, despite noting a trade-off between empathy and coherence. Liu et al., conversely, integrated emotional and visual modalities using Emotional Retrieval Module(ERM), Response Emotion Prediction (REP), and Emotion-Enhanced Response Generation (EERG) models. Their multimodal ELMD system demonstrated superior emotional and contextual responsiveness compared to traditional dialogue systems.
@@
Chen et al. Chen2024 , Motwani et al. Motwani2024 , and Yang et al. Yang2025 further highlighted emotional considerations, albeit within specialised frameworks. Chen et al.’s RECONCILE framework collaboratively utilised multiple LLM agents to generate emotionally and contextually relevant responses, leveraging weighted voting based on reasoning justification. Motwani et al. proposed a sequential training method employing specialised agents in reasoning and verification tasks, indirectly enhancing emotional intelligence through robust error correction. Yang et al.’s SupportlyChat directly focused on therapeutic contexts using Cognitive Behavioural Therapy (CBT), combining RoBERTa and sentiment analysis to enhance emotional awareness and professional interaction in mental health support scenarios.
@@
Finally, Brun et al. Brun2025 evaluated emotional sensitivity in technical support chatbots, specifically employing VADER sentiment analysis within ChatGPT-3.5. Their findings indicated emotional sensitivity significantly improved user-perceived competence, empathy, trust, and continued use intention, although without measurably affecting user emotional states.
@@ 2.2 Multi-Agent Debate and Collaborative Reasoning
Research exploring multi-agent systems frequently emphasises collaborative reasoning and decision-making through debate mechanisms. Zhao et al. Zhao2024 tested electoral methods from social choice theory (e.g., Borda, IRV, Ranked Pairs) in collective decision-making scenarios involving multiple LLM agents, highlighting notable accuracy improvements in benchmarks like MMLU and ARC. They demonstrated that even minimal agent ensembles significantly enhance performance and recommended task-specific tailoring of decision methods.
Yang et al. Yang2025a combined adversarial debate, weighted voting, and internal self-correction within a multi-agent framework to mitigate hallucinations and enhance reasoning efficiency. Their methodology involved iterative debate, error documentation, and dynamic weighting, outperforming baseline methods in accuracy and response speed on tasks like GSM8K and MMLU.
Similarly, Xu et al. Xu2023 proposed a peer-review inspired collaborative method. Agents independently generated solutions, reviewed peers’ contributions, and refined responses accordingly. This explicit feedback and diverse model integration consistently improved performance across multiple datasets (GSM8K, StrategyQA, ARC-c).
@@
In contrast to existing approaches, the architecture proposed in this work distinguishes itself by employing a structured multi-agent framework in which each agent represents a distinct emotion and participates in iterative emotional reasoning. While prior studies tend to integrate emotion either through static embeddings, sentiment lexicons, or personalised memory modules Ratican2023 ; Kang2024 ; Rasool2024 , this architecture fosters dynamic emotional dialogue and collective deliberation. Additionally, the integration of symbolic reasoning mechanisms—such as voting and justification—enhances explainability, a feature largely absent in previous neural architectures Chen2024 ; Motwani2024 . The system is also extensible to different emotional models and has been adapted for high-stakes domains, such as emergency response, through the Armando variant. To the best of our knowledge, no previous work combines emotional simulation, multimodal reasoning, and factual grounding within a unified architecture.
@@ 3 Proposed Architecture
This paper presents an innovative conversational AI system architecture that processes user queries through distinct emotional lenses, specifically a set of five basic emotions: Anger, Joy, Sadness, Fear, and Disgust. The system follows the emotional framework portrayed in the first Inside Out film, which was inspired by Paul Ekman’s theory of universal emotions Ekman1971 . Although Ekman initially identified six basic emotions—Happiness, Anger, Sadness, Disgust, Surprise, and Fear—and later expanded the list to include emotions such as Contempt and Enjoyment Ekman , the film deliberately reduced the set to five for narrative clarity Cannon2016 ; Roper2015 . In this work, the same five emotions were retained to preserve the analogy with the film, thereby fostering greater familiarity and intuitive use—particularly among users who have seen the film. However, it is important to emphasise that the Project Riley architecture is not limited to this configuration: it is capable of simulating a broader range of emotions, depending on the requirements of the application and the computational resources available.
The system can be easily adapted to reflect alternative emotional models, allowing developers to incorporate different sets of emotions based on specific application needs or cultural relevance. Furthermore, the proposed architecture can be aligned with other psychological theories, such as Plutchik’s Wheel of Emotions Plutchik1982 , Carroll Izard’s Differential Emotions Theory Izard1977 , Jaak Panksepp’s Affective Neuroscience framework Panksepp1982 , or dimensional approaches like Russell’s Circumplex Model Russell1980 and the Pleasure-Arousal-Dominance (PAD) model Mehrabian1980 . Such adaptations would require architectural modifications and prompt reengineering to ensure consistency with the theoretical assumptions of the chosen model.
@@ 4 Project Riley chatbot prototype
In accordance with the proposed architecture, a general-purpose chatbot was implemented based entirely on the Project Riley architecture. To enable local deployment of LLM, the Ollama framework was utilised Ollama2025 . The hardware specifications of the machine used for this implementation are detailed in Table 1.
@@
Given the hardware constraints, particularly regarding Graphics Processing Unit (GPU) Video Random Access Memory (VRAM) capacity, considerable attention was directed towards selecting models that offer a balanced trade-off between performance and model size. Furthermore, when computationally feasible, we prioritised the use of abliterated models. Abliteration is a technique that alters the weights of a model to bypass default safety alignment mechanisms typically introduced during fine-tuning Labonne2024 . These mechanisms, while intended to ensure safe output, often over-filter emotionally charged or controversial prompts, leading to overly polite, neutral, or evasive responses.
@@
[…]
@@ 4.1 Prototype Evaluation
Analysis of responses to the question Which emotion do you predominantly identify in the final response? (LABEL:tab:predominant_emotions) reveals significant variability in emotional perception. In the Anxiety in Academic or Professional Contexts theme, Joy was most frequently identified (7 mentions), possibly reflecting a reassuring approach adopted by the system. Conversely, Sadness dominated responses related to Breakups or Friendship Loss (5 mentions), indicating the appropriateness of a more reflective emotional tone in relational loss scenarios. Difficult Personal Decisions equally elicited Fear and Joy (5 mentions each), suggesting responses were balanced between cautionary and optimistic perspectives. Family or Intergenerational Conflicts predominantly evoked Joy (6 mentions), again highlighting positivity in familial contexts. Lastly, Job Loss or Unemployment mainly elicited Fear (5 mentions), aligning well with the typically uncertain and stressful nature of employment-related concerns.
@@ 5 Armando: A emergency response chatbot using Project Riley
This approach seeks to mitigate issues such as those observed on 28 April 2025, during a widespread power outage across the Iberian Peninsula, when some individuals turned to ChatGPT for information. As the chatbot was not updated with official communications from public safety authorities, its use led to the dissemination of unreliable and inaccurate information, thereby increasing public confusion and anxiety SAPO242025 .
[…]
@@
Given the specific requirements of this emergency response application, namely, reduced response times and lower computational overhead-targeted optimisations were implemented. The Final Synthesis and Voting and Analysis processes were simplified to enhance execution speed and reduce GPU VRAM usage. To this end, prompt modifications were introduced to emulate a reasoning logic using a conventional text-based LLM. Furthermore, as the THOUGHTS section was deemed non-essential in an emergency context and its exclusion contributed to faster response generation, it was omitted from the chatbot’s Final Synthesis.
@@
As illustrated in Figure 5, RAG enrichment is applied immediately before the final synthesis stage, contrary to what might appear more intuitive, which would be to introduce it prior to the multi-round processing. Such an approach would provide each emotional agent with validated information from the outset, allowing the entire discussion to be grounded in authoritative content. However, initial development and empirical testing revealed that this strategy does not reliably ensure the factual accuracy of the final output, as critical information may become diluted or distorted during the emotional dialogue. Therefore, RAG enrichment was instead positioned just before the synthesis stage, ensuring that the final response maintains emotional balance and contextual sensitivity, while also conveying accurate and up-to-date information to the user.